OpenStudy (anonymous):

Felicia's doctor has put her on a diet of 500 calories per meal. This means she can consume no more than 500 calories at any meal. Her options for lunch in the cafeteria are as follows: Yogurt that has 110 calories per serving Mini Sandwiches that have 100 calories per serving Which inequality below can be used to find the possible combinations of servings of yogurt (y) and mini sandwiches (s) she can eat and stay within her goal of no more than 500 calories?

OpenStudy (anonymous):

110y + 100s >= 500 110y - 100s >= 500 110y + 100s <= 500 110y - 100s <= 500

OpenStudy (anonymous):

The options are all very similar. So look at the difference between the answers and try to reason it out a bit. If Felicia is going to eat more than one thing, will her calories add up or subtract?

OpenStudy (anonymous):

add so would it be C?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

yw